# Decentralized Zeno-Free Event-Triggered Control For Multiple Networks Subject to Stochastic Network Delays and Poisson Pulsing Attacks

### v1

@@ II-B Stochastic Hybrid Formalism
where $x\in\mathbb{R}^{n}$ denotes the system state, $F$ and $G$ denote flow and jump set-valued mappings, respectively, the flow set $C$ and the jump set $D$ respectively allow the flowing and jumping states. Notation $x^{+}$ denotes the instantaneous change depending on jumping states. Placeholder $v^{+}$ holds the place for a sequence of independent, identically distributed (i.i.d.) input random variables $\{\textbf{v}_{n}\}_{n=1}^{\infty}$, which is defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, i.e., $\textbf{v}_{n}:\Omega\rightarrow\mathbb{R}^{m}$ for $n\in\mathbb{Z}_{\geq 1}$. Given the Borel $\sigma$-field $\textbf{B}(\mathbb{R}^{m})$ over $\mathbb{R}^{m}$, it holds that $\textbf{v}^{-1}_{n}(A):=\{\omega\in\Omega:\textbf{v}_{n}(\omega)\in A\}\in\mathcal{F}$, and $\mathbb{P}\{\omega\in\Omega:\textbf{v}_{n}(\omega)\in A\}$ is well-defined and independent of $n$, for $A\in\textbf{B}(\mathbb{R}^{m})$. Notation $\mu$ denotes the distribution function derived from $(\Omega,\mathcal{F},\mathbb{P})$, then $\mu:\textbf{B}(\mathbb{R}^{m})\rightarrow[0,1]$ holds with $\mu(A):=\mathbb{P}\{\omega\in\Omega:\textbf{v}_{n}(\omega)\in A\}$, see [15, Sec. 2.1 and 11.1]. Denote the formalism of the stochastic hybrid system (II-B) as $\mathcal{H}:=(C,F,D,G,\mu)$, where the solution’ uniqueness is not assumed due to set-valued flow/jump mappings and the potentially non-trivial overlap of flow and jump sets, see [1].
@@
The following mild regularity conditions are adopted for the stochastic hybrid system (II-B) from [1, Assum. 1-2].
@@
Explanations on these regularity conditions in Assumption 1: Random solutions to (II-B) are ensured simultaneously by the local boundedness (see items A2) and 3a)) and by the measurability (see item 3b)) [16, Sec. IV]; and additionally, the nominal robustness is guaranteed by the local boundedness and the outer semicontinuity, which enables the equivalence between non-uniform and uniform versions of stochastic stability properties, see [17, Claim 1, Thms. 2-3] and [18, Thms. 4-5] for the difference inclusions, as well as [19] for (II-B). The measurability in item 3b) implies that $v\mapsto G(x,v)$ is measurable for any $x\in\mathbb{R}^{n}$, see [20, Prop. 2], [21, Remark 1] and [14, Thm. 14.13(b)], used to establish certain well-posed integrals used later; and additionally, the condition of closed values in item 3b) is equivalent to the outer semicontinuity of $x\mapsto G(x,v)$ for each $v\in\mathbb{R}^{n}$, see [14, Thm. 5.7(a)]. As analyzed in [1] and [22], item 3b) holds if $G(x,v)$ is outer semicontinuous; holds if the domain of $v\mapsto G(x,v)$ is countable; also holds if a single-valued mapping $G$ is a Carathéodory mapping [14, Exam. 14.15], i.e., $x\mapsto G(x,v)$ is continuous and $v\mapsto G(x,v)$ is measurable. We note that Assumption 1 essentially agrees with the non-stochastic hybrid basic assumptions in [23, Assumption 6.5], when $G$ does not depend on $v$. A relationship route is presented in Fig. 1 for Assumption 1-3b).
@@ II-C Networked Control Configuration
This paper utilizes stochastic hybrid analysis tools to facilitate more efficient analysis and ensure more reliable conclusions. The networked control system under consideration, with state $x_{p}\in\mathbb{R}^{n_{p}}$, is represented as the stochastic hybrid formalism (II-B) through the emulation method [12]: A controller with state $x_{c}\in\mathbb{R}^{n_{c}}$ is first designed in advance that stabilizes the system in the absence of multiple networks that operate asynchronously and independently (i.e., decentralized), then implemented over the attacked networks to maintain the stability by designing the decentralized triggered strategies which generate transmission instants $\{t^{i}_{j}\}_{j\in\mathbb{Z}_{\geq 0}}$ satisfying $0\leq t^{i}_{j}<t^{i}_{j+1}$ for the $i$th-network, $i\in\bar{N}$. In what follows, two sources of randomness are considered in multiple decentralized networks: the networked communication is constrained by the stochastic network delays and also by Poisson pulsing denial-of-service (Pp-DoS) attacks, where “pulsing” here means each attack pulse lasts for a negligible short period of time (each attack’s duration is negligible). The network delay denotes the continuous time from a transmission instant (sampled data enters the network) to the corresponding update instant (sampled data leaves the network), and is supposed to be subject to certain continuous probability distribution; while the attacks’ cardinal number is a discrete random variable supposed to be subject to a Poisson distribution. We assume there is no communication delay or sampling delay outside the networks.
@@
Figure 3: A typical evolution of $x_{pi\ell}$ and $\hat{x}_{pi\ell}$ for $\ell=2$, where $t^{i}_{j-1}$, $t^{i}_{j}$ and $t^{i}_{j+1}$ are three consecutive transmission instants for $x_{pi}=(x_{pi1},x_{pi2})$ in the absence of attacks. Notations $\bullet$ and $\circ$ represent that the piecewise constant $\hat{x}_{pi\ell}$ is left-continuous. (a) No attacks in $[t^{i}_{j},t^{i}_{j}+\tau^{i}_{j}]$. (b) A Pp-DoS attack occurs at update instant $t^{i}_{j}+\tau^{i}_{j}$. See explanations above the model (II-C).
@@
where notation $u\in\mathbb{R}^{n_{u}}$ denotes the control input. Functions $f_{p}$ and $f_{c}$ are supposed to be continuous; and (supposed to be) continuously differentiable function $g_{c}$ can be relaxed to be discontinuous, since Assumption 1 can also be ensured by getting the convex closure of discontinuous functions, see [23, pg. 74-75]. Denote $\mathcal{E}$ as the closed equilibrium point set of the closed-loop $(x_{p},x_{c})$-system (II-C).
@@
Figure 4: Evolution illustrations for $x_{pi}$ and $\hat{x}_{pi}$, where $t^{i}_{j-1}$, $t^{i}_{j}$ and $t^{i}_{j+1}$ are three consecutive transmission instants for $x_{pi}$ in the absence of attacks. Notations $\bullet$ and $\circ$ represent that the piecewise constant $\hat{x}_{pi}$ is left-continuous. (a) No attacks in $[t^{i}_{j},t^{i}_{j}+\tau^{i}_{j}]$. (b) A Pp-DoS attack occurs at update instant $t^{i}_{j}+\tau^{i}_{j}$. See explanations below the system model (II-C).
@@ III Mathematical Formulation
Before further analyzing the stability property of system (II-C) using stochastic hybrid analysis tools, we need to model (II-C) as the stochastic hybrid formalism (II-B). To do so, we first define the following notations: $x:=(\tilde{x},e,\tau_{e},k,s,l,m,\chi)\in\mathbb{X}$, $\tilde{x}:=(x_{p},x_{c})\in\mathbb{R}^{n_{\tilde{x}}}$, $\mathbb{X}:=\mathbb{R}^{n_{\tilde{x}}}\times\mathbb{R}^{n_{p}}\times\mathbb{R}^{N}_{\geq 0}\times\mathbb{Z}^{N}_{\geq 0}\times\mathbb{R}^{n_{p}}\times\{0,1\}^{N}\times\{0,1\}^{N}\times\mathbb{R}^{N}_{\geq 0}$, $\tau_{e}:=(\tau^{1}_{e},\ldots,\tau^{N}_{e})\in\mathbb{R}^{N}_{\geq 0}$ with $\tau^{i}_{e}\in\mathbb{R}_{\geq 0}$ capturing the time elapsed since the latest transmissions, $k:=(k_{1},\ldots,k_{N})\in\mathbb{Z}^{N}_{\geq 0}$ with $k_{i}\in\mathbb{Z}_{\geq 0}$ keeping track of the total amount of transmissions over time, $s:=(s_{1},\ldots,s_{N})\in\mathbb{R}^{n_{p}}$ with $s_{i}\in\mathbb{R}^{n_{p,i}}$ serving as a memory variable to store the value $h_{i}(j,e_{i}(t^{i}_{j}))-e_{i}(t^{i}_{j})$ at the moment of a transmission at time $t^{i}_{j}$ and is used to model the update event at time $t^{i}_{j}+\tau^{i}_{j}$ (see (6)), $l:=(l_{1},\ldots,l_{N})\in\{0,1\}^{N}$ with $l_{i}=0$ ($l_{i}=1$) keeping track of whether the next event is transmission (or update), $m:=(m_{1},\ldots,m_{N})\in\{0,1\}^{N}$ with $m_{i}=0$ ($m_{i}=1$) indicating whether the attack occurs or not (below (8)) and thus keeping track of whether the next update is successful (or not), $\chi:=(\chi_{1},\ldots,\chi_{N})\in\mathbb{R}^{N}_{\geq 0}$ with $\chi_{i}$ denoting the triggered function (see (9)) for the $i$th network $\mathcal{N}_{i}$, $i\in\bar{N}$. Combining (II-C), (4), (7) and (9) implies the flow map in (II-B), i.e.,
@@
with the jump set $C:=\bigcap_{i\in\bar{N}}C_{i}$, $C_{i}:=\{x\in\mathbb{X}|((\tau^{i}_{e}\leq\tau^{m_{i},i}_{miet}~{}\vee~{}\chi_{i}\geq 0)\wedge(l_{i}=0))\vee((0\leq\tau^{i}_{e}\leq\tau^{i}_{mad})\wedge(l_{i}=1))\}$; and also implies the jump map in (II-B), i.e., $G(x,v^{+})=\bigcup_{i\in\bar{N}}G_{i}(x,v^{+})$ and
@@
for $h(k,e):=(h_{i}(k_{i},e_{i}))_{i\in\bar{N}}$ where $h_{i}(k_{i},e_{i})$ corresponds to the term $h_{i}(j,e_{i})$ in (7), implying $G_{0,0,i}$ corresponds to a successful transmission, $G_{1,0,i}$ corresponds to a successful update, while $G_{1,1,i}$ corresponds to a failed update. That is, the goal of this paper is to make the closed set $\mathcal{A}:=\{x\in\mathbb{X}|\tilde{x}=\mathcal{E},e=s=\textbf{0}_{n_{p}},\chi=\textbf{0}_{N}\}$ uniformly globally asymptotically stable in probability for the SHS (II-B) with (10) and (11).
This paper assumes that almost each stochastic hybrid trajectory starts in the set $\mathbb{X}_{0}=\{x\in\mathbb{X}|\tau_{e}\geq\tau^{0,i}_{miet},s=0,\chi=0\}$. The initial set can be freely chosen and is deterministic once chosen, which does not influences the global sense: No matter how far the solutions start from $\mathcal{A}$ within a distance $\delta>0$, the probability $\rho$ of having solutions escaping $\mathcal{A}+\varepsilon\mathbb{B}^{\circ}$ can be arbitrarily small for certain $\varepsilon>0$ (see definition below). We note that choosing $\mathbb{X}_{0}$ does not impose any constraints on initial states $x_{p}$ and $x_{c}$, or initial networked value $\hat{x}_{p}$ at the controller side, but rather reflects the initialization of variables in the triggered strategy (8)-(9). In the following, we report the concept for the uniform global asymptotic stability in probability (UGASp) of a closed set for SHS (II-B), which is with respect to the probability where the graphs of solutions simultaneously have three subproperties, i.e., uniform global stability in probability (UGSp) composed of uniform Lyapunov stability in probability (ULySp) and uniform Lagrange stability in probability (ULaSp), and uniform global attractivity in probability (UGAp). To do so, we define $\Gamma_{\geq\tau}:=\{(s,t)\in\mathbb{R}^{2}:s+t\geq\tau\}$, $\mathcal{S}_{r}(K)$ as the set of solutions to (II-B) with initial conditions belonging to $K$, and $\mbox{graph}(\textbf{x}(\omega)):=\{(t,j,z)\in\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n}:(t,j)\in\mbox{dom}~{}\textbf{x}_{\omega},z=\textbf{x}_{\omega}(t,j),\textbf{x}(\omega):=\textbf{x}_{\omega}\}$, where $\omega$ can be dropped on a few occasions without confusions.

### v3

# A Stochastic Hybrid Approach to Decentralized Networked Control: Stochastic Network Delays and Poisson Pulsing Attacks

@@ II-B Stochastic Hybrid Formalism
where $x\in\mathbb{R}^{n}$ denotes the system state, $F$ and $G$ denote flow and jump set-valued mappings, respectively, the flow set $C$ and the jump set $D$ respectively allow the flowing and jumping states. Notation $x^{+}$ denotes the instantaneous change depending on jumping states. Placeholder $v^{+}$ holds the place for a sequence of independent, identically distributed (i.i.d.) input random variables $\{\textbf{v}_{n}\}_{n=1}^{\infty}$, which is defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, i.e., $\textbf{v}_{n}:\Omega\rightarrow\mathbb{R}^{m}$ for $n\in\mathbb{Z}_{\geq 1}$. Given the Borel $\sigma$-field $\textbf{B}(\mathbb{R}^{m})$ over $\mathbb{R}^{m}$, it holds that $\textbf{v}^{-1}_{n}(A):=\{\omega\in\Omega:\textbf{v}_{n}(\omega)\in A\}\in\mathcal{F}$, and $\mathbb{P}\{\omega\in\Omega:\textbf{v}_{n}(\omega)\in A\}$ is well-defined and independent of $n$, for $A\in\textbf{B}(\mathbb{R}^{m})$. Notation $\mu$ denotes the distribution function derived from $(\Omega,\mathcal{F},\mathbb{P})$, then $\mu:\textbf{B}(\mathbb{R}^{m})\rightarrow[0,1]$ holds with $\mu(A):=\mathbb{P}\{\omega\in\Omega:\textbf{v}_{n}(\omega)\in A\}$, see [15, Sec. 2.1 and 11.1]. Denote the formalism of the stochastic hybrid system (1) as $\mathcal{H}:=(C,F,D,G,\mu)$, where the solution’ uniqueness is not assumed due to set-valued flow/jump mappings and the potentially non-trivial overlap of flow and jump sets, see [1].
@@
The following mild regularity conditions are adopted for the stochastic hybrid system (1) from [1, Assum. 1-2].
@@
Explanations on these regularity conditions in Assumption 1: Random solutions to (1) are ensured simultaneously by the local boundedness (see items A2) and 3a)) and by the measurability (see item 3b)) [16, Sec. IV]; and additionally, the nominal robustness is guaranteed by the local boundedness and the outer semicontinuity, which enables the equivalence between non-uniform and uniform versions of stochastic stability properties, see [17, Claim 1, Thms. 2-3] and [18, Thms. 4-5] for the difference inclusions, as well as [19] for (1). The measurability in item 3b) implies that $v\mapsto G(x,v)$ is measurable for any $x\in\mathbb{R}^{n}$, see [20, Prop. 2], [21, Remark 1] and [14, Thm. 14.13(b)], used to establish certain well-posed integrals used later; and additionally, the condition of closed values in item 3b) is equivalent to the outer semicontinuity of $x\mapsto G(x,v)$ for each $v\in\mathbb{R}^{n}$, see [14, Thm. 5.7(a)]. As analyzed in [1] and [22], item 3b) holds if $G(x,v)$ is outer semicontinuous; holds if the domain of $v\mapsto G(x,v)$ is countable; also holds if a single-valued mapping $G$ is a Carathéodory mapping [14, Exam. 14.15], i.e., $x\mapsto G(x,v)$ is continuous and $v\mapsto G(x,v)$ is measurable. We note that Assumption 1 essentially agrees with the non-stochastic hybrid basic assumptions in [23, Assumption 6.5], when $G$ does not depend on $v$. A relationship route is presented in Fig. 1 for Assumption 1-3b).
@@ II-C Networked Control Configuration
This paper utilizes stochastic hybrid analysis tools to facilitate more efficient analysis and ensure more reliable conclusions. The networked control system under consideration, with state $x_{p}\in\mathbb{R}^{n_{p}}$, is represented as the stochastic hybrid formalism (1) through the emulation method [12]: A controller with state $x_{c}\in\mathbb{R}^{n_{c}}$ is first designed in advance that stabilizes the system in the absence of multiple networks that operate asynchronously and independently (i.e., decentralized), then implemented over the attacked networks to maintain the stability by designing the decentralized triggered strategies which generate transmission instants $\{t^{i}_{j}\}_{j\in\mathbb{Z}_{\geq 0}}$ satisfying $0\leq t^{i}_{j}<t^{i}_{j+1}$ for the $i$th-network, $i\in\bar{N}$. In what follows, two sources of randomness are considered in multiple decentralized networks: the networked communication is constrained by the stochastic network delays and also by Poisson pulsing denial-of-service (Pp-DoS) attacks, where “pulsing” here means each attack pulse lasts for a negligible short period of time (each attack’s duration is negligible). The network delay denotes the continuous time from a transmission instant (sampled data enters the network) to the corresponding update instant (sampled data leaves the network), and is supposed to be subject to certain continuous probability distribution; while the attacks’ cardinal number is a discrete random variable supposed to be subject to a Poisson distribution. We assume there is no communication delay or sampling delay outside the networks.
@@
Figure 3: A typical evolution of $x_{pi\ell}$ and $\hat{x}_{pi\ell}$ for $\ell=2$, where $t^{i}_{j-1}$, $t^{i}_{j}$ and $t^{i}_{j+1}$ are three consecutive transmission instants for $x_{pi}=(x_{pi1},x_{pi2})$ in the absence of attacks. Notations $\bullet$ and $\circ$ represent that the piecewise constant $\hat{x}_{pi\ell}$ is left-continuous. (a) No attacks in $[t^{i}_{j},t^{i}_{j}+\tau^{i}_{j}]$. (b) A Pp-DoS attack occurs at update instant $t^{i}_{j}+\tau^{i}_{j}$. See explanations above the model (2).
@@
where notation $u\in\mathbb{R}^{n_{u}}$ denotes the control input. Functions $f_{p}$ and $f_{c}$ are supposed to be continuous; and (supposed to be) continuously differentiable function $g_{c}$ can be relaxed to be discontinuous, since Assumption 1 can also be ensured by getting the convex closure of discontinuous functions, see [23, pg. 74-75]. Denote $\mathcal{E}$ as the closed equilibrium point set of the closed-loop $(x_{p},x_{c})$-system (2).
@@
Figure 4: Evolution illustrations for $x_{pi}$ and $\hat{x}_{pi}$, where $t^{i}_{j-1}$, $t^{i}_{j}$ and $t^{i}_{j+1}$ are three consecutive transmission instants for $x_{pi}$ in the absence of attacks. Notations $\bullet$ and $\circ$ represent that the piecewise constant $\hat{x}_{pi}$ is left-continuous. (a) No attacks in $[t^{i}_{j},t^{i}_{j}+\tau^{i}_{j}]$. (b) A Pp-DoS attack occurs at update instant $t^{i}_{j}+\tau^{i}_{j}$. See explanations below the system model (2).
@@ III Mathematical Formulation
Before further analyzing the stability property of system (2) using stochastic hybrid analysis tools, we need to model (2) as the stochastic hybrid formalism (1). To do so, we first define the following notations: $x:=(\tilde{x},e,\tau_{e},k,s,l,m,\chi)\in\mathbb{X}$, $\tilde{x}:=(x_{p},x_{c})\in\mathbb{R}^{n_{\tilde{x}}}$, $\mathbb{X}:=\mathbb{R}^{n_{\tilde{x}}}\times\mathbb{R}^{n_{p}}\times\mathbb{R}^{N}_{\geq 0}\times\mathbb{Z}^{N}_{\geq 0}\times\mathbb{R}^{n_{p}}\times\{0,1\}^{N}\times\{0,1\}^{N}\times\mathbb{R}^{N}_{\geq 0}$, $\tau_{e}:=(\tau^{1}_{e},\ldots,\tau^{N}_{e})\in\mathbb{R}^{N}_{\geq 0}$ with $\tau^{i}_{e}\in\mathbb{R}_{\geq 0}$ capturing the time elapsed since the latest transmissions, $k:=(k_{1},\ldots,k_{N})\in\mathbb{Z}^{N}_{\geq 0}$ with $k_{i}\in\mathbb{Z}_{\geq 0}$ keeping track of the total amount of transmissions over time, $s:=(s_{1},\ldots,s_{N})\in\mathbb{R}^{n_{p}}$ with $s_{i}\in\mathbb{R}^{n_{p,i}}$ serving as a memory variable to store the value $h_{i}(j,e_{i}(t^{i}_{j}))-e_{i}(t^{i}_{j})$ at the moment of a transmission at time $t^{i}_{j}$ and is used to model the update event at time $t^{i}_{j}+\tau^{i}_{j}$ (see (6)), $l:=(l_{1},\ldots,l_{N})\in\{0,1\}^{N}$ with $l_{i}=0$ ($l_{i}=1$) keeping track of whether the next event is transmission (or update), $m:=(m_{1},\ldots,m_{N})\in\{0,1\}^{N}$ with $m_{i}=0$ ($m_{i}=1$) indicating whether the attack occurs or not (below (8)) and thus keeping track of whether the next update is successful (or not), $\chi:=(\chi_{1},\ldots,\chi_{N})\in\mathbb{R}^{N}_{\geq 0}$ with $\chi_{i}$ denoting the triggered function (see (9)) for the $i$th network $\mathcal{N}_{i}$, $i\in\bar{N}$. Combining (2), (4), (7) and (9) implies the flow map in (1), i.e.,
@@
with the jump set $C:=\bigcap_{i\in\bar{N}}C_{i}$, $C_{i}:=\{x\in\mathbb{X}|((\tau^{i}_{e}\leq\tau^{m_{i},i}_{miet}~{}\vee~{}\chi_{i}\geq 0)\wedge(l_{i}=0))\vee((0\leq\tau^{i}_{e}\leq\tau^{i}_{mad})\wedge(l_{i}=1))\}$; and also implies the jump map in (1), i.e., $G(x,v^{+})=\bigcup_{i\in\bar{N}}G_{i}(x,v^{+})$ and
@@
for $h(k,e):=(h_{i}(k_{i},e_{i}))_{i\in\bar{N}}$ where $h_{i}(k_{i},e_{i})$ corresponds to the term $h_{i}(j,e_{i})$ in (7), implying $G_{0,0,i}$ corresponds to a successful transmission, $G_{1,0,i}$ corresponds to a successful update, while $G_{1,1,i}$ corresponds to a failed update. That is, the goal of this paper is to make the closed set $\mathcal{A}:=\{x\in\mathbb{X}|\tilde{x}=\mathcal{E},e=s=\textbf{0}_{n_{p}},\chi=\textbf{0}_{N}\}$ uniformly globally asymptotically stable in probability for the SHS (1) with (10) and (11).
This paper assumes that almost each stochastic hybrid trajectory starts in the set $\mathbb{X}_{0}=\{x\in\mathbb{X}|\tau_{e}\geq\tau^{0,i}_{miet},s=0,\chi=0\}$. The initial set can be freely chosen and is deterministic once chosen, which does not influences the global sense: No matter how far the solutions start from $\mathcal{A}$ within a distance $\delta>0$, the probability $\rho$ of having solutions escaping $\mathcal{A}+\varepsilon\mathbb{B}^{\circ}$ can be arbitrarily small for certain $\varepsilon>0$ (see definition below). We note that choosing $\mathbb{X}_{0}$ does not impose any constraints on initial states $x_{p}$ and $x_{c}$, or initial networked value $\hat{x}_{p}$ at the controller side, but rather reflects the initialization of variables in the triggered strategy (8)-(9). In the following, we report the concept for the uniform global asymptotic stability in probability (UGASp) of a closed set for SHS (1), which is with respect to the probability where the graphs of solutions simultaneously have three subproperties, i.e., uniform global stability in probability (UGSp) composed of uniform Lyapunov stability in probability (ULySp) and uniform Lagrange stability in probability (ULaSp), and uniform global attractivity in probability (UGAp). To do so, we define $\Gamma_{\geq\tau}:=\{(s,t)\in\mathbb{R}^{2}:s+t\geq\tau\}$, $\mathcal{S}_{r}(K)$ as the set of solutions to (1) with initial conditions belonging to $K$, and $\mbox{graph}(\textbf{x}(\omega)):=\{(t,j,z)\in\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n}:(t,j)\in\mbox{dom}~{}\textbf{x}_{\omega},z=\textbf{x}_{\omega}(t,j),\textbf{x}(\omega):=\textbf{x}_{\omega}\}$, where $\omega$ can be dropped on a few occasions without confusions.